NEED TO SOLVE.
A.𝑓(𝑥)=x2+2x; 𝑓(1),𝑓(−2),𝑓(12),𝑓(𝑎),𝑓(−𝑎),𝑓(𝑎+𝑏)
B.𝑔(𝑥)=1−x1+x; 𝑔(2),𝑔(−2),𝑔(12),𝑔(𝑎),𝑔(𝑎−1),𝑔(−1)
C.𝑓(𝑥)=22+3𝑥−4; 𝑓(0),𝑓(2),𝑓(−2),𝑓(√2),𝑓(𝑥+1),𝑓(−𝑥)
D.𝑓(𝑥)=𝑥3−4𝑥2; 𝑓(0),𝑓(1),𝑓(−1),𝑓(32),𝑓(𝑥2),𝑓(𝑥2)
E.𝑓(𝑥)=2𝑥𝑥−1
F.𝑓(𝑥)=3−5𝑥+4𝑥2
G.𝑓(𝑥)=1𝑥−3
H.𝑓(𝑥)=𝑥+2𝑥2−1
I.𝑓(𝑥)=√𝑥−5
J.ℎ(𝑥)=√2𝑥−5

THANK YOU!

Answer by Alan3354(69443)

NEED TO SOLVE.
A.��(��)=x2+2x; ��(1),��(−2),��(12),��(��),��(−��),��(��+��)
B.��(��)=1−x1+x; ��(2),��(−2),��(12),��(��),��(��−1),��(−1)
C.��(��)=22+3��−4; ��(0),��(2),��(−2),��(√2),��(��+1),��(−��)
D.��(��)=��3−4��2; ��(0),��(1),��(−1),��(32),��(��2),��(��2)
E.��(��)=2����−1
F.��(��)=3−5��+4��2
G.��(��)=1��−3
H.��(��)=��+2��2−1
I.��(��)=√��−5
J.ℎ(��)=√2��−5
-----------------
I don't know what all the rectangles mean.